Eukaryotic DNA goes through multiple "levels" of packing. Let's assume that each level of packing reduces the length of a piece of DNA by 12-fold (so the resulting DNA is 1/12 its previous length after each fold)/ The largest chromosome in this species is 0.8 billion base pairs long. Based on this information, how many "levels" of packing would be needed in order to fit the DNA into a nucleus 1.5 micrometers in diameter?
Answer #1

Ans. Packing ratio of DNA= Length of DNA/ the length into which it is packaged

Nucleus is in the circular form then first of all find out nucleus length by applying formula 2πr

Diameter of nucleus is = 1.5 × 10-6 micrometers.

Radius of nucleus = 1.5 × 10-6 /2 ( Radius= Diameter/2)

= 0.75/ 106

= 0.00000075

Radius of nucleus = 0.00000075 micrometers

Length (circumference) of a nucleus= 2πr

= 2× 3.14 × 0.00000075

=0.00000471

Packing ratio= length of DNA/ length of nucleus

= 0.8 × 109 / 0.00000471

= 0.8 × 109/ 0.00000471

= 169851.4 × 109